OpenStudy (anonymous):

Find the simple interest earned after 12 years on an investment of $5,000 at 4% interest earned annually.

OpenStudy (mven):

5000*0.04=X then X*12=V V= the interest after 12 years the total cost you can find by adding V to 5000

OpenStudy (anonymous):

So $2400? c:

OpenStudy (mven):

oops im sorry i meant 0.04*12=x then add that to 5000

OpenStudy (mven):

so $5000.48 i believe